# Ali Dasdan KD Consulting Saratoga, CA, USA

### v2

@@ Abstract
Applied to productivity competition, we prove firms face a coordination failure preventing productivity gains from translating fully to worker welfare. We show this formally for working hours; the same structural logic extends to wages, benefits, and job security. European evidence confirms that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, productivity welfare decoupling persists, manifesting differently across institutional contexts.
@@
The analysis is diagnostic rather than prescriptive, identifying structural barriers to coordination rather than proposing solutions whose feasibility our framework questions.
@@ 1.1 The Productivity Puzzle
John Maynard Keynes famously predicted in 1930 that technological progress would reduce the working week to 15 hours by century’s end (Keynes, 1930). Nearly a century later, this prediction appears dramatically wrong. Workers globally typically work 1,350-2,200+ hours annually (OECD, 2023), far exceeding Keynes’s predicted 780 hours, and in many sectors working hours have increased rather than decreased (Schor, 1992); this competitive escalation is exemplified by the “996” work culture (9 am to 9 pm, 6 days a week) in China (Wikipedia contributors, 2025), which has even been emulated by some US startups seeking to match that intensity (The New York Times, 2025).
@@
This study is also motivated by a persistent empirical pattern observed by the author in the software and technology sectors over the last two decades. In these industries, leadership consistently demands output expansion following productivity gains, often reducing engineering headcount rather than stabilizing it. This ’do more with less’ dynamic persists even when cost-cutting threatens organizational sustainability.
@@ 1.4 Preview of Contributions
Third, we apply this framework to productivity competition and show that it satisfies all five conditions, explaining Keynes’s failed prediction structurally rather than culturally. We formalize the coordination problem firms face when productivity increases, showing why welfare improvements remain difficult despite technological capacity. European evidence shows that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, coordination achieves only partial, costly overrides of the tragic logic that erode over time.
@@ 2.1 The Prisoner’s Dilemma and Social Dilemmas
Schelling (1960, 1973) extended this analysis to N-player settings with binary choices and externalities. When multiple actors [^1] We use ’player’ and ’actor’ interchangeably for game participants. face similar decisions where individual choices impose externalities on others, coordination failures can emerge even when all actors understand the collective problem. The key insight is that as $N$ increases, the individual incentive to deviate grows stronger—each actor’s restraint provides smaller individual benefit while the temptation to defect remains constant.
@@
Jervis (1978) applied this logic to international relations, analyzing the “security dilemma” (originally coined by Herz, 1950) where states’ efforts to increase their own security—through arms buildups or territorial expansion—threaten other states and trigger countermeasures that leave all parties less secure. The security dilemma exhibits the core social dilemma structure: each state has incentive to arm regardless of others’ choices (dominance), yet mutual restraint would benefit all (Pareto-efficiency), but international anarchy prevents enforceable agreements. Jervis demonstrated that this tragic dynamic persists even when all states prefer peace, because the anarchic structure of international relations makes credible commitment to cooperation impossible.
@@ 2.3 Productivity and Working Hours
Recent work by Acemoglu and Restrepo (2020) and Acemoglu (2025) provides important insights into technology adoption, distinguishing between automation that displaces labor (“so-so automation”) and technology that creates new tasks or augments workers (“human-augmenting”). They argue that market forces may lead to socially excessive automation due to insufficient incentives for firms to internalize benefits that accrue to workers.
While Schnaiberg (1980) intuited competitive pressure as a “treadmill,” and Acemoglu and Restrepo (2020); Acemoglu (2025) identified market failures in technology choices, no formal game-theoretic model explains the coordination problem firms face when productivity increases. Also while extensive research documents the secular decline in the labor share of income globally, (Karabarbounis and Neiman, 2014; Elsby et al., 2013), the literature lacks a unified game-theoretic explanation for why competitive pressure systematically prevents productivity gains from translating to worker welfare.
We provide the first formal game-theoretic model of productivity competition as an N-player social dilemma. Building on Schnaiberg (1980)’s treadmill intuition, we prove that firms face a structural coordination failure: each firm has a dominant strategy to expand output rather than reduce hours, yet universal restraint would Pareto-dominate universal expansion. We formalize the barriers to coordination that make this tragedy persist. This structural framework explains Schnaiberg (1980)’s empirical observation, integrates with Acemoglu and Restrepo (2020); Acemoglu (2025)’s analysis of technology choice failures, and clarifies why the regulatory interventions documented by Alesina et al. (2005) require costly, persistent institutional support to partially override market dynamics.
@@ 2.4 AI Governance and Safety
Bostrom (2014) provided a comprehensive analysis of existential risks from superintelligent AI, emphasizing the orthogonality thesis (intelligence and goals are independent) and instrumental convergence (advanced AI systems will pursue convergent subgoals regardless of final objectives). Russell (2019) framed AI safety as an inverse reinforcement learning problem and proposed provably beneficial AI through value alignment. Amodei et al. (2016) identified concrete problems in AI safety: avoiding negative side effects, reward hacking, distributional shift, safe exploration, and robustness to distributional change. Dafoe (2018) analyzed AI governance challenges, identifying the need for international coordination on development and deployment. Bengio et al. (2025) provided a comprehensive synthesis of the current evidence on the capabilities, risks, and safety of advanced AI systems. Panigrahy and Sharan (2025) proved that under reasonable mathematical definitions of safety, trust, and artificial general intelligence (AGI), a safe and trusted AI system cannot be an AGI system.
Armstrong et al. (2016) analyzed “racing to the precipice”, the concern that competitive pressure could drive unsafe AI development. Their analysis identifies the race dynamic but does not provide a formal structural model explaining why coordination fails.
@@ 3 A Unified Framework for Structural Tragedies
We use the term “Tragedy” in the strict sense defined by Whitehead (1925) and popularized by Hardin (1968): “the solemnity of the remorseless working of things.” A situation becomes a tragedy not through error or malice, but through inevitability. Specifically, it occurs when individual rationality compels defection (C3) even though actors collectively recognize cooperation as superior (C4), while structural barriers prevent coordination (C5). The tragedy lies in this structural trap: actors are not making mistakes—they are imprisoned by the logic of their own survival.
We formalize the payoff structure as follows. Let $s=(s_{1},s_{2},\ldots,s_{n})$ be a strategy profile, which is a vector specifying the chosen strategy $s_{i}\in\{C,D\}$ for every player $i$.
@@ 3.1 The Five Conditions
A problem exhibits structural tragedy if and only if it satisfies five conditions:
@@
The binary structure simplifies the game while capturing the essential strategic tension. In practice, choices may appear continuous, but the strategic logic often reduces to a binary decision: restrain (cooperate) or expand (defect). The key is that individual choices impose externalities—one player’s defection makes cooperation more costly for others.
@@
The equilibrium outcome is recognizably suboptimal—all players would prefer mutual cooperation to mutual defection if they could coordinate.
This condition ensures the tragedy is recognizable. Players understand that mutual defection makes everyone worse off than mutual cooperation. The tragedy is not hidden—it is visible, predictable, and collectively unwanted. Yet the structure makes it difficult to escape.
@@
This condition distinguishes structural tragedies from simple coordination problems. Many Pareto-inefficient equilibria can be overcome through contracts, monitoring, or institutional design. Structural tragedies persist because enforcement mechanisms face fundamental barriers—not merely high costs but structural impossibilities.
Throughout the paper, we will refer to the conditions using the condition number and the first letter of the condition name as in “C1:N” or “C5:E”.
@@ 3.2 The Formal Result
We now prove these conditions are both necessary and sufficient to generate structural tragedy.
@@
We establish sufficiency and necessity separately.
@@
Claim (i): Unique Nash Equilibrium
@@
Claim (iii): Pareto-Domination
@@
We prove necessity by showing the contrapositive: if any of Conditions 3-5 fails to hold, then the tragedy structure breaks (at least one of claims (i)-(iv) fails). By contrapositive logic, this establishes that each condition is necessary for tragedy.
@@
  • Actions are perfectly observable and verifiable, or
  • Monitoring is feasible at reasonable cost, or
@@
Under these conditions, players can contract to implement $s^{C}$ and sustain it through enforcement mechanisms. Defection can be detected and punished. Commitments become credible. The folk theorem for repeated games may apply: if players are sufficiently patient and monitoring is adequate, cooperation can be sustained as a subgame perfect equilibrium through trigger strategies that punish defection.
While cooperation may require institutional infrastructure, it becomes achievable through standard mechanisms—contracts, monitoring, reputation, repeated interaction. The tragedy dissolves from structurally impossible to merely institutionally challenging. Therefore C5 is necessary.
We have shown that removing any single condition eliminates the structural tragedy (contrapositive of necessity). Combined with Part I showing sufficiency, we conclude that Conditions 1-5 are necessary and sufficient for structural tragedy as characterized by Theorem 6.
This theorem establishes that Conditions 1-5 precisely characterize structural tragedies. The equilibrium is not merely Pareto-inefficient: it is recognizably suboptimal yet structurally difficult to escape. The framework provides both diagnostic power (checking conditions predicts tragedy) and intervention guidance (breaking any condition resolves tragedy).
@@ 3.3 Validation Across Canonical Cases
We now validate this framework by applying it to established coordination failures across multiple domains. Table 1 summarizes the analysis.
Table 1: Five Conditions Across Canonical Coordination Failures, including the two new ones introduced in this paper.
Domain C1:N C2:B C3:D C4:P C5:E Reference
@@
We briefly discuss each case, demonstrating how it satisfies all five conditions.
@@
Commons Tragedies (Hardin, 1968): Multiple shepherds (N-player) each decide whether to add sheep to common pasture (binary choice creating negative externality through overgrazing). Each shepherd benefits from adding sheep regardless of others’ choices (dominance driven by individual profit maximization). Yet universal restraint would preserve the commons for all (Pareto-domination). Absent property rights or enforceable agreements (enforcement difficulty), the commons is depleted.
@@
Security Dilemmas (Jervis, 1978; Mearsheimer, 2001): Multiple states (N-player) each decide whether to arm or remain unarmed (binary choice where one state’s arms create security threats for others). Each state benefits from arming for security regardless of others’ choices (dominance driven by survival concerns in anarchy). Yet universal disarmament would provide security without arms costs (Pareto-domination). International anarchy prevents enforceable disarmament agreements (enforcement difficulty).
@@
Bank Runs (Diamond and Dybvig, 1983): Multiple depositors (N-player) each decide whether to withdraw or keep deposits in a bank (binary choice where withdrawals create panic). Each depositor benefits from withdrawing if others might withdraw (dominance driven by fear of losing deposits). Yet if no one panics, all maintain deposits safely (Pareto-domination). Absence of instant coordination during panics (enforcement difficulty) allows runs to occur. Deposit insurance (FDIC) broke this tragedy (for deposits up to the insurance limit) by eliminating existential stakes, removing Condition 3. Note that the wholesale funding and equities markets are not covered under the FDIC insurance so they may still be vulnerable to bank runs.
@@
Space Debris (Adilov et al., 2015): Multiple states and firms (N-player) each decide whether to limit launches or expand operations (binary choice where debris from launches threatens all satellites). Each actor benefits from launches regardless of others’ choices (dominance driven by commercial and security advantages). Yet universal restraint would preserve orbital space (Pareto-domination). International space law lacks enforcement mechanisms (enforcement difficulty).
@@
Antibiotic Resistance (Foster and Grundmann, 2006): Multiple actors including doctors, patients, farmers (N-player) each decide whether to limit or freely use antibiotics (binary choice where overuse breeds resistance affecting all). Each actor benefits from use regardless of others’ choices (dominance driven by immediate health or productivity gains). Yet universal restraint would preserve antibiotic effectiveness (Pareto-domination). Distributed usage across multiple sectors resists coordination (enforcement difficulty).
@@
Public Goods (Olson, 1965): Multiple individuals (N-player) each decide whether to contribute to a public good (Cooperate) or withhold contribution (Defect). Withholding allows an individual to save resources (free-ride) while still enjoying any good provided by others (externality). Each individual has a dominant incentive to free-ride, as their personal contribution is costly and has minimal impact on the total outcome (dominance). Yet, universal contribution would fund the public good, making everyone better off than if the good is not provided (Pareto-domination). Absent a coercive mechanism like taxation (enforcement difficulty), the public good is under-provided or not provided at all.
@@
Hockey Helmets (Schelling, 1973): Multiple players (N-player) each decide whether to wear a helmet (Cooperate/Restrain) or play without one (Defect/Advantage-seeking). Playing without a helmet provides a slight competitive advantage (e.g., better vision, “toughness”) but increases the overall risk of injury for all (negative externality). Each player benefits from the competitive edge of not wearing one, regardless of others’ choices (dominance). Yet, universal helmet use would neutralize the competitive effects and make everyone safer (Pareto-domination). Absent a league-wide, binding mandate (enforcement difficulty), players are incentivized to play without helmets despite the collective danger.
@@
Climate Change (Ostrom, 2010): Multiple nations and firms (N-player) each decide whether to reduce emissions (Cooperate) or pollute freely (Defect). Polluting provides immediate economic advantages while imposing global climate damage (negative externality). Each actor benefits from polluting to maintain economic competitiveness, regardless of others’ choices (dominance). Yet, universal restraint would prevent catastrophic climate costs, making all better off (Pareto-domination). International anarchy and sovereignty concerns prevent binding, enforceable agreements (enforcement difficulty).
@@
Result: Perfect verification across all canonical cases. Each tragedy satisfies the complete set of conditions. This suggests the framework captures the essential structure that creates coordination failures.

### v3

@@ Abstract
Applied to productivity competition, we prove firms face a coordination failure preventing productivity gains from translating fully to worker welfare. We show this formally for working hours; the same structural logic extends to wages, benefits, and job security. European evidence confirms that, even under maximally favorable institutional conditions, productivity welfare decoupling persists, manifesting differently across institutional contexts.
@@
The analysis is diagnostic rather than prescriptive: we synthesize from the literature structural barriers both to standard solutions and to the fundamental changes that would be required to resolve these tragedies, rather than proposing implementable solutions whose feasibility our framework questions.
@@ 1.1 The Productivity Puzzle
John Maynard Keynes famously predicted in 1930 that technological progress would reduce the working week to 15 hours by century’s end [^1] While Keynes’s phrasing was speculative (“Three-hour shifts or a fifteen-hour week may put off the problem”), the literature has standardly interpreted this as a prediction of future working hours (e.g., (Crafts, 2022)). (Keynes, 1930). Nearly a century later, this prediction appears dramatically wrong. Workers globally typically work 1,350-2,200+ hours annually (OECD, 2025), far exceeding Keynes’s predicted 780 hours, and in many sectors working hours have increased rather than decreased (Schor, 1992); this competitive escalation is exemplified by the ‘996’ work culture (9 am to 9 pm, 6 days a week) in China (Wikipedia contributors, 2025), which has even been emulated by some US startups seeking to match that intensity (The New York Times, 2025).
@@
This study is also motivated by a persistent empirical pattern observed by the author in the software and technology sectors over the last two decades. In these industries, leadership consistently demands output expansion following productivity gains, often reducing engineering headcount rather than stabilizing it. This ‘do more with less’ dynamic persists even when cost-cutting threatens organizational sustainability.
@@ 1.4 Preview of Contributions
Third, we apply this framework to productivity competition and show that it satisfies all five conditions, explaining Keynes’s failed prediction structurally rather than culturally. We formalize the coordination problem firms face when productivity increases, showing why welfare improvements remain difficult despite technological capacity. European evidence shows that, even under maximally favorable institutional conditions such as wealthy democracies, strong institutions, and supranational governance, coordination achieves only partial, costly overrides of the tragic logic that erode over time.
@@ 2.1 The Prisoner’s Dilemma and Social Dilemmas
Schelling (1960, 1973) extended this analysis to N-player settings with binary choices and externalities. When multiple actors [^2] We use ‘player’ and ‘actor’ interchangeably for game participants. face similar decisions where individual choices impose externalities on others, coordination failures can emerge even when all actors understand the collective problem. The key insight is that as N increases, the individual incentive to deviate grows stronger—each actor’s restraint provides smaller individual benefit while the temptation to defect remains constant.
@@
Jervis (1978) applied this logic to international relations, analyzing the “security dilemma” (originally coined by Herz (1950)) where states’ efforts to increase their own security—through arms buildups or territorial expansion—threaten other states and trigger countermeasures that leave all parties less secure. The security dilemma exhibits the core social dilemma structure: each state has incentive to arm regardless of others’ choices (dominance), yet mutual restraint would benefit all (Pareto-efficiency), but international anarchy prevents enforceable agreements. Jervis demonstrated that this tragic dynamic persists even when all states prefer peace, because the anarchic structure of international relations makes credible commitment to cooperation impossible.
@@ 2.3 Productivity and Working Hours
Recent work by Acemoglu and Restrepo (2020) and Acemoglu (2025) provides important insights into technology adoption, distinguishing between automation that displaces labor (“so-so automation”) and technology that creates new tasks or augments workers (“human-augmenting”). They argue that market distortions lead to socially excessive investment in displacement.
While these theories explain specific drivers—Schnaiberg (1980) on competitive intuition and Acemoglu and Restrepo (2020); Acemoglu (2025) on technology choice—the literature lacks a unified game-theoretic explanation for the coordination problem in allocation. We distinguish between which technologies firms adopt (Acemoglu’s focus) and how productivity gains are utilized (our focus). Also, while extensive research documents the secular decline in the labor share of income globally (Karabarbounis and Neiman, 2014; Elsby et al., 2013), a formal model explaining why competitive pressure systematically prevents even human-augmenting productivity gains from translating to worker welfare remains absent.
We provide the first formal game-theoretic model of productivity competition as an N-player social dilemma. Building on Schnaiberg (1980)’s treadmill intuition, we prove that firms face a structural coordination failure: each firm has a dominant strategy to expand output rather than reduce hours, yet universal restraint would Pareto-dominate universal expansion. We formalize the barriers to coordination that make this tragedy persist. This structural framework explains Schnaiberg (1980)’s empirical observation, complements Acemoglu and Restrepo (2020); Acemoglu (2025)’s analysis of adoption failures with a theory of allocation failures, and clarifies why the regulatory interventions documented by Alesina et al. (2005) require costly, persistent institutional support to partially override market dynamics.
@@ 2.4 AI Governance and Safety
Bostrom (2014) provided a comprehensive analysis of existential risks from superintelligent AI, emphasizing the orthogonality thesis (intelligence and goals are independent) and instrumental convergence (advanced AI systems will pursue convergent subgoals regardless of final objectives). Russell (2019) framed AI safety as an inverse reinforcement learning problem and proposed provably beneficial AI through value alignment. Amodei et al. (2016) identified concrete problems in AI safety: avoiding negative side effects, reward hacking, distributional shift, safe exploration, and robustness to distributional change. Dafoe (2018) analyzed AI governance challenges, identifying the need for international coordination on development and deployment. Bengio et al. (2025) provided a comprehensive synthesis of the current evidence on the capabilities, risks, and safety of advanced AI systems. Panigrahy and Sharan (2025) proved that under reasonable mathematical definitions of safety, trust, and artificial general intelligence (AGI), a safe and trusted AI system cannot be an AGI system. Armstrong et al. (2016) analyzed “racing to the precipice”, the concern that competitive pressure could drive unsafe AI development. Their analysis identifies the race dynamic but does not provide a formal structural model explaining why coordination fails.
@@ 3 A Unified Framework for Structural Tragedies
We use the term “Tragedy” in the strict sense defined by Whitehead (1925) and popularized by Hardin (1968): “the solemnity of the remorseless working of things.” A situation becomes a tragedy not through error or malice, but through inevitability. Specifically, it occurs when individual rationality compels defection even though actors collectively recognize cooperation as superior, while structural barriers prevent coordination. The tragedy lies in this structural trap: actors are not making mistakes—they are imprisoned by the logic of their own survival.
We formalize the payoff structure as follows. Let $s=(s_{1},s_{2},\ldots,s_{n})$ be a strategy profile for nn=N players, which is a vector specifying the chosen strategy $s_{i}\in\{C,D\}$ for every player $i$, where $C$ represents Cooperate (restraint) and $D$ represents Defect (advantage-seeking).
@@ 3.1 The Five Conditions
A problem exhibits structural tragedy if and only if it satisfies the following five conditions:
@@
The binary structure simplifies the game while capturing the essential strategic tension. In practice, choices may appear continuous, but the strategic logic often reduces to a binary decision: cooperate or defect. The key is that individual choices impose externalities: one player’s defection makes cooperation more costly for others.
@@
The equilibrium outcome is recognizably suboptimal: all players would prefer mutual cooperation to mutual defection if they could coordinate.
This condition ensures the tragedy is recognizable. Players understand that mutual defection makes everyone worse off than mutual cooperation. The tragedy is not hidden; it is visible, predictable, and collectively unwanted. Yet the structure makes it difficult to escape.
@@
This condition distinguishes structural tragedies from simple coordination problems. Many Pareto-inefficient equilibria can be overcome through contracts, monitoring, or institutional design. Structural tragedies persist because enforcement mechanisms face fundamental barriers: not merely high costs but severe structural obstacles.
@@ 3.2 The Formal Result
We now prove these five conditions are both necessary and sufficient to generate structural tragedy.
@@
We establish sufficiency and necessity separately, followed by the overall conclusion.
@@
Claim (i): Unique Nash Equilibrium.
@@
Claim (iii): Pareto-Domination.
@@
We prove necessity by showing the contrapositive: if any of Conditions 1-5 fails to hold, then the tragedy structure breaks (at least one of claims (i)-(iv) fails). By contrapositive logic, this establishes that each condition is necessary for tragedy.
@@
If there is a single decision-maker ($n=1$) or a central authority with binding power over all players, then there is no coordination problem. The single player or central authority simply implements the Pareto-optimal outcome $s^{C}$ directly. No strategic interaction exists, and no tragedy arises. Therefore C1 is necessary.
@@
If individual choices create no externalities, each player’s payoff depends only on their own choice, not others’ choices: $U_{i}(s_{i},s_{-i})=U_{i}(s_{i})$. Each player independently optimizes their own choice without strategic consideration of others’ actions. There is no coordination problem because choices are independent. Strategic interaction disappears, eliminating the coordination failure. Therefore C2 is necessary.
@@
If defection is not strictly dominant, then there exist some player $i$ and some opponent profile $s_{-i}$ such that $U_{i}(C,s_{-i})\geq U_{i}(D,s_{-i})$. In such cases, cooperation can be a best response to others cooperating. This creates the possibility of multiple Nash equilibria, including $s^{C}$ as a stable equilibrium. The tragedy is no longer structurally inevitable. Therefore C3 is necessary.
@@
If the equilibrium $s^{D}$ is not Pareto-dominated by $s^{C}$, then for some player $i$ we have $U_{i}(s^{D})\geq U_{i}(s^{C})$, or universal defection is Pareto-optimal. Then the equilibrium outcome is efficient by definition: there exists no alternative outcome that makes everyone better off. Without Pareto-domination, there is no tragedy. The outcome may be a coordination problem, but it is not recognizably suboptimal. Therefore C4 is necessary.
@@
Suppose structural barriers to enforcement are absent. Specifically, suppose any of the following holds:
@@
  • Actions are sufficiently observable that defection can be reliably detected, or
  • Monitoring is feasible at reasonable cost relative to coordination benefits, or
@@
Under these conditions, the folk theorem for repeated games applies: cooperation can be sustained as a subgame perfect equilibrium through trigger strategies because defection is detectable and punishable. For the structural tragedy to hold, C5 must therefore imply barriers sufficient to preclude these solutions: specifically, monitoring must be too imperfect to detect defection, or patience too low for future consequences to matter. Absent such barriers, the problem dissolves from a structural tragedy to a solvable coordination challenge. Therefore C5 is necessary.
Conclusion.
[…]
This theorem formally distinguishes structural tragedies from solvable coordination problems. By establishing that the tragedy persists if and only if all five conditions hold, the result confirms that the equilibrium is structurally resilient against standard interventions. This provides precise diagnostic utility: any proposed solution that fails to break at least one condition is mathematically guaranteed to fail, while effective governance requires targeting the specific structural barriers identified by the framework.
@@ 3.3 Validation Across Canonical Cases
We now validate this framework by applying it to established coordination failures across multiple domains. Table 1 summarizes the analysis. We briefly discuss each case, demonstrating how it satisfies all five conditions.
Table 1: Five Conditions C1-C5 Across Canonical Coordination Failures, including the two new ones introduced in this paper.
Domain C1 C2 C3 C4 C5 Reference
@@
Commons Tragedies
@@
(Hardin, 1968): Multiple shepherds (N-player) each decide whether to add sheep to common pasture (binary choice creating negative externality through overgrazing). Each shepherd benefits from adding sheep regardless of others’ choices (dominance driven by individual profit maximization). Yet universal restraint would preserve the commons for all (Pareto-domination). Absent property rights or enforceable agreements (enforcement difficulty), the commons is depleted.
Security Dilemmas
@@
(Jervis, 1978; Mearsheimer, 2001): Multiple states (N-player) each decide whether to arm or remain unarmed (binary choice where one state’s arms create security threats for others). Each state benefits from arming for security regardless of others’ choices (dominance driven by survival concerns in anarchy). Yet universal disarmament would provide security without arms costs (Pareto-domination). International anarchy prevents enforceable disarmament agreements (enforcement difficulty).
Bank Runs
@@
(Diamond and Dybvig, 1983): Multiple depositors (N-player) each decide whether to withdraw or keep deposits in a bank (binary choice where withdrawals create panic). Each depositor benefits from withdrawing if others might withdraw (dominance driven by fear of losing deposits). Yet if no one panics, all maintain deposits safely (Pareto-domination). Absence of instant coordination during panics (enforcement difficulty) allows runs to occur. Deposit insurance (FDIC) broke this tragedy (for deposits up to the insurance limit) by eliminating existential stakes, removing Condition 3. Note that the wholesale funding and equities markets are not covered under the FDIC insurance so they may still be vulnerable to bank runs.
Space Debris
@@
(Adilov et al., 2015): Multiple states and firms (N-player) each decide whether to limit launches or expand operations (binary choice where debris from launches threatens all satellites). Each actor benefits from launches regardless of others’ choices (dominance driven by commercial and security advantages). Yet universal restraint would preserve orbital space (Pareto-domination). International space law lacks enforcement mechanisms (enforcement difficulty).
Antibiotic Resistance
@@
(Foster and Grundmann, 2006): Multiple actors including doctors, patients, farmers (N-player) each decide whether to limit or freely use antibiotics (binary choice where overuse breeds resistance affecting all). Each actor benefits from use regardless of others’ choices (dominance driven by immediate health or productivity gains). Yet universal restraint would preserve antibiotic effectiveness (Pareto-domination). Distributed usage across multiple sectors resists coordination (enforcement difficulty).
Public Goods
@@
(Olson, 1965): Multiple individuals (N-player) each decide whether to contribute to a public good (Cooperate) or withhold contribution (Defect). Withholding allows an individual to save resources (free-ride) while still enjoying any good provided by others (externality). Each individual has a dominant incentive to free-ride, as their personal contribution is costly and has minimal impact on the total outcome (dominance). Yet, universal contribution would fund the public good, making everyone better off than if the good is not provided (Pareto-domination). Absent a coercive mechanism like taxation (enforcement difficulty), the public good is under-provided or not provided at all.
Hockey Helmets
@@
(Schelling, 1973): Multiple players (N-player) each decide whether to wear a helmet (Cooperate/Restrain) or play without one (Defect/Advantage-seeking). Playing without a helmet provides a slight competitive advantage (e.g., better vision, ‘toughness’) but increases the overall risk of injury for all (negative externality). Each player benefits from the competitive edge of not wearing one, regardless of others’ choices (dominance). Yet, universal helmet use would neutralize the competitive effects and make everyone safer (Pareto-domination). Absent a league-wide, binding mandate (enforcement difficulty), players are incentivized to play without helmets despite the collective danger.
Climate Change
@@
(Ostrom, 2010): Multiple nations and firms (N-player) each decide whether to reduce emissions (Cooperate) or pollute freely (Defect). Polluting provides immediate economic advantages while imposing global climate damage (negative externality). Each actor benefits from polluting to maintain economic competitiveness, regardless of others’ choices (dominance). Yet, universal restraint would prevent catastrophic climate costs, making all better off (Pareto-domination). International anarchy and sovereignty concerns prevent binding, enforceable agreements (enforcement difficulty).
Result:
@@
Perfect verification across all canonical cases. Each tragedy satisfies the complete set of conditions. This suggests the framework captures the essential structure that creates coordination failures.